OpenStudy (anonymous):

Help please1. (–10x3 + 30x – 20) ÷ (–5x + 5) (1 point ) A.2x2 – 2x + 4 B.–2x2 – 2x – 4 C.–2x2 + 2x + 4 D.2x2 + 2x – 4

Parth (parthkohli):

The last one is the answer. You get the answer by factoring. (-5x + 5)(2x2 + 2x - 4) = (-10x3 + 30x - 20)

OpenStudy (anonymous):

so it's d?

Parth (parthkohli):

Yes.

OpenStudy (anonymous):

1. (–10x3 + 30x – 20) ÷ (–5x + 5) (1 point) (0 pts) 2x2 – 2x + 4 (0 pts) –2x2 – 2x – 4 (0 pts) –2x2 + 2x + 4 (1 pt) 2x2 + 2x – 4 1 /1 point 2. The width w of a rectangular swimming pool is x + 4. The area A of the pool is 2x3 – 29x + 12. What is an expression for the length of the pool? (1 point) (0 pts) 2x2 + 8x +3 (0 pts) 2x2 – 8x – 3 (1 pt) 2x2 – 8x + 3 (0 pts) 2x2 + 8x – 3 0 /1 point 3. Katalin drove 180 miles on her vacation. She drove an average of 1.5 times faster on the second 90 miles of her trip than she did on the first 90 miles of her trip. Which expression represents the time she spent driving? Let x = her speed on the first half of the trip. (1 point) (1 pt) (0 pts) (0 pts) 150x (0 pts) 1 /1 point 4. What is the graph of ? (1 point) (0 pts) (0 pts) (0 pts) (1 pt) 0 /1 point 5. A software company is releasing one of their products on a CD. The manufacturer charges a $5,500 setup fee and $1.00 per CD. Approximately how many CDs must be produced so that the total cost per CD is $3.50? (1 point) (0 pts) 1,222 CDs (1 pt) 2,200 CDs (0 pts) 1,571 CDs (0 pts) 5,500 CDs 1 /1 point 6. Simplify. (1 point) (0 pts) (1 pt) (0 pts) (0 pts) 1 /1 point 7. Simplify. (1 point) (0 pts) (0 pts) (0 pts) (1 pt) 1 /1 point 8. Simplify. (1 point) (1 pt) (0 pts) (0 pts) (0 pts) x – 7 1 /1 point 9. Simplify. (1 point) (1 pt) x – 5; where x –2 (0 pts) x – 5; where x 5 (0 pts) (0 pts) x – 2 1 /1 point 10. Simplify. (1 point) (0 pts) (0 pts) (1 pt) (0 pts) 1 /1 point 11. Multiply. (1 point) (0 pts) (0 pts) (1 pt) (k + 3)(3k + 2) (0 pts) (3k + 1)(2k + 3) 1 /1 point 12. Divide. (1 point) (0 pts) (0 pts) (0 pts) (1 pt) 0 /1 point 13. Divide. (1 point) (1 pt) (0 pts) (0 pts) (0 pts) 0 /1 point 14. Divide. (8m7 – 10m5) ÷ 2m3 (1 point) (0 pts) 4m7 – 5m5 (0 pts) 4m4 – 10m5 (0 pts) 8m7 – 10m2 (1 pt) 4m4 –5m2 1 /1 point 15. Divide. (–6m9 – 6m8 – 16m6) ÷ (2m3) (1 point) (0 pts) –3m9 – 3m8 – 8m6 (0 pts) –3m6 – 6m8 – 16m3 (1 pt) –3m6 – 3m5 – 8m3 (0 pts) –3m6 – 3m5 – 16m3 1 /1 point 16. Add or subtract. (1 point) (0 pts) (0 pts) (1 pt) (0 pts) 0 /1 point 17. Add or subtract. (1 point) (1 pt) (0 pts) (0 pts) (0 pts) 0 /1 point 18. Add or subtract. (1 point) (0 pts) (1 pt) (0 pts) (0 pts) 1 /1 point 19. Add or subtract. (1 point) (0 pts) (1 pt) (0 pts) (0 pts) 0 /1 point 20. Solve. (1 point) (0 pts) x = – (0 pts) x = – (0 pts) x = (1 pt) x = –2 1 /1 point 21. Solve. (1 point) (0 pts) x = 1 (1 pt) (0 pts) x = – (0 pts) x = –5 1 /1 point 22. Solve. (1 point) (0 pts) x = 5, –3 (1 pt) x = –5, –3 (0 pts) x = –5, 3 (0 pts) x = 5, 3 0 /1 point 23. Bob and James are finishing the roof of a house. Working alone, Bob can shingle the roof in 10 hours. James can shingle the same roof in 16 hours. How long will it take them working together to shingle the roof? Round your answer to the nearest hundredth if necessary. (1 point) (0 pts) 13 hours (0 pts) 26.67 hours (1 pt) 6.15 hours (0 pts) 0.04 hours 1 /1 point 24. The temperature T of a given mass of gas varies inversely with its volume V. The temperature of 90 cm3 of a certain gas is 30°C. What will be the temperature of the gas when it is compressed to a volume of 18 cm3? (1 point) (0 pts) 6°C (1 pt) 150°C (0 pts) 120°C (0 pts) 75°C 0 /1 point 25. The time t required to drive a certain distance varies inversely with the speed r. If it takes 4 hours to drive the distance at 35 miles per hour, how long will it take to drive the same distance at 45 miles per hour? (1 point) (1 pt) about 3.11 hours (0 pts) 140 hours (0 pts) about 5.14 hours (0 pts) 393.75 hours 1 /1 point 26. Do the data in the table represent a direct variation or an inverse variation? Write an equation to model the data in the table. (1 point) (0 pts) direct variation; y = x (0 pts) inverse variation; xy = 7 (1 pt) direct variation; y = 7x (0 pts) inverse variation; xy = 1 /1 point 27. What are the excluded values of the function? (1 point) (0 pts) 0.04 (0 pts) 24 (1 pt) –24 (0 pts) –0.04 1 /1 point 28. Identify the asymptotes of the graph. (1 point) (1 pt) (0 pts) (0 pts) (0 pts) 0 /1 point